# Uncertainties from metrology in the integrated luminosity measurement with the updated design of the very forward region of a detector at CEPC

### v1

@@ 4 Conclusion
Although the method of integrated luminosity has already been studied in great detail at LEP, proposals for new $e^{+}e^{-}$ colliders call for quantification of achievable luminosity precision in each individual case. We do it here for CEPC, from the perspective of mechanical and beam-related requirements at the $Z^{0}$ pole, with detector placed on the z-axis.
At the $Z^{0}$ pole, control of the luminometer inner radius at the micrometer level is posing the most demanding requirement regarding detector manufacturing, regardless of the luminometer’s positioning (on the s-axis or on the z-axis). It is important to note that eventual changes in the luminometer design towards smaller polar angle coverage will require control of the luminometer inner radius even below micrometer precision, according to the dependence of the Bhabha cross-section on the polar angles of the scattered particles. However, these results are obtained under assumption that the change of the luminometer’s inner radius corresponds to exactly the same change of the inner radius of the luminometer’s fiducial volume, which needs to be approved. Also, control of the asymmetrical bias in beam energies might be challenging, while the beam energy spread can be relaxed significantly beyond the current half-width of 18 MeV at the $Z^{0}$ pole (0.04%, Table 1).
From the point of view of experimental input for the Bhabha cross-section calculations, center-of-mass energy should be known at the $\mathrm{10^{-4}}$ level, which requires further studies in terms of feasibility of such a precision. This might be an issue for the cross-section measurements of the slope of a production threshold.
Although the simulated Bhabha samples are statistically small to be exact on boundaries of the considered systematics, distribution of central values presented in this study confirms that considering the mechanics and positioning, as well as the beam-related uncertainties, no effect seems to be more critical for precision measurement of the integrated luminosity if the luminometer is placed around the z-axis than if it is placed around the outgoing beams. The LEP-style asymmetric counting, however, will not be effective for the luminometer placed on the z-axis to cope with the effects affecting the left-right symmetry.
Complex metrology, together with the beam-related interactions affecting the Bhabha final state at small polar angles may call for consideration of an alternative central process, like di-muon or di-photon production for the integrated luminosity measurement, with hopefully less complex systematic effects subjected to further studies. In addition, studies presented here should be complemented with a full simulation study of systematic effects arising from the detector design, technology and performance in the presence of realistic backgrounds.

### v2

@@ 4 Conclusion
Although the method of integrated luminosity has been studied in a great detail at LEP (LEP1 , LEP2 ), for each newly proposed $e^{+}e^{-}$ collider it is necessary to quantify the achievable luminosity precision, in particular at the $Z^{0}$ resonance. Here it is done for CEPC, taking into consideration the mechanical and beam-related requirements at the $Z^{0}$ pole, with the luminometer placed on the z-axis. It is the first attempt at a future Higgs factory to quantify systematic effects rising from metrology, with luminometer displaced from the the outgoing beams where it is conventionally positioned at colliders with a crossing angle.
Control of the luminometer inner radius at the micrometer level at the $Z^{0}$ pole seeems to be the most demanding requirement for the detector manufacturing, regardless of the luminometer’s positioning (on the s-axis or on the z-axis). Any eventual change in the luminometer design towards smaller polar angle coverage will require control of the luminometer inner radius even below micrometer precision, according to the dependence of the Bhabha cross-section on the polar angles of the scattered particles. However, it is important to note that these results are obtained under assumption that the change of the luminometer’s inner radius corresponds to exactly the same change of the inner radius of the luminometer’s fiducial volume, which needs to be approved. Also, control of the asymmetrical bias in beam energies might be challenging, while the beam energy spread can be relaxed significantly beyond the current RMS of 18 MeV at the $Z^{0}$ pole.
From the point of view of experimental input for the Bhabha cross-section calculations, center-of-mass energy should be known at the $\mathrm{10^{-4}}$ level, which requires further studies in terms of feasibility of such a precision. This might be an issue for the cross-section measurements on the slope of a production threshold.
Although the simulated Bhabha samples are relatively small to be exact on boundaries of the considered systematics, distribution of central values presented in this study confirms that no effect seems to be more critical for precision measurement of the integrated luminosity if the luminometer is placed around the z-axis than if it is placed around the outgoing beams. The LEP-style asymmetric counting, however, will not be effective for the luminometer placed on the z-axis to cope with the effects arising from the left-right symmetry.
Complex metrology, together with the beam-related interactions affecting the Bhabha final state at small polar angles Voutsinas may call for consideration of an alternative or a complementary central process, like di-muon or di-photon production for the integrated luminosity measurement, with hopefully less complex systematic effects subjected to further studies ECFA . In addition, studies presented here should be complemented with a full simulation study of systematic effects arising from the detector design, technology and performance in the presence of realistic backgrounds.